# Large Language Models are Contrastive Reasoners

### v1

@@ Abstract
Prompting methods play a crucial role in enhancing the capabilities of pre-trained large language models (LLMs). We explore how contrastive prompting (CP) significantly improves the ability of large language models to perform complex reasoning. We demonstrate that LLMs are decent contrastive reasoners by simply adding “Let’s give a correct and a wrong answer.” before LLMs provide answers. Experiments on two large language models show that zero-shot contrastive prompting improves performance on a range of arithmetic, commonsense, and symbolic reasoning tasks without any hand-crafted few-shot examples, such as increasing the accuracy on GSM8K from $35.9\%$ to $88.8\%$ and AQUA-RAT from $41.3\%$ to $62.2\%$ with the state-of-the-art GPT-4 model. Our method not only surpasses zero-shot CoT and few-shot CoT in most arithmetic and commonsense reasoning tasks but also can seamlessly integrate with existing prompting methods, resulting in improved or comparable results when compared to state-of-the-art methods. Our code is available at the following GitHub repository: https://github.com/yao8839836/cp.
@@ 1 Introduction
Recent studies [27, 2, 12] have shown that large language models (LLMs) exhibit impressive performance across a wide range of tasks. In particular, the chain-of-thought (CoT) prompting technique has demonstrated the capability of LLMs to handle complex tasks, including math problem solving, by guiding them to generate intermediate reasoning steps [21, 9, 26]. These studies spotlight the significance of developing efficient techniques to direct LLMs in their reasoning processes [10, 1, 4, 24].
Nevertheless, the current chain-of-thought (CoT) paradigm encounters two main challenges: offering correct guidance or examples of reasoning and reducing the reliance on manual labeling. In particular, zero-shot CoT [9] provides general reasoning guidance by providing instructions like “Think step by step.”, but the generated reasoning steps may not be correct and adequate for tasks such as commonsense question-answering (Section 4). On the other hand, few-shot CoT [21] offers more detailed guidance but necessitates labeled examples of the reasoning process, which can be expensive to obtain for each task. This raises an important research question: Is it possible to generate a more accurate reasoning process without relying on human labeling?
In this work, we introduce contrastive prompting, a novel prompting approach that automatically directs the reasoning process of large language models. Our inspiration stems from how humans can learn from both their correct and incorrect actions [15]. For instance, when confronted with a math problem (as in Figure 1), people may ask "How can we prevent mistakes in each step?" By identifying the steps that are prone to mistakes on their own, they can enhance their ability to avoid mistakes and provide accurate solutions. Our idea is to prompt LLMs to emulate this reasoning process, enabling them to effectively solve new problems.
Specifically, when presented with a problem to solve, we instruct LLMs to generate both correct and incorrect answers within the given context. To achieve this, we provide prompts such as "Let’s give a correct and a wrong answer." Following this, we verify and confirm the correct answer. Our proposed approach offers multiple advantages. It not only generates incorrect answers autonomously but also places a greater emphasis on ensuring the accuracy of the answers. This eliminates the need for manually labeling reasoning examples for each task and problem, effectively addressing the challenges faced by zero-shot and few-shot CoT.
We evaluate the proposed approach across various reasoning-intensive tasks, including arithmetic reasoning (SingleEq, AddSub, MultiArith, AQUA-RAT, GSM8K, SVAMP), commonsense reasoning (CommonsenseQA, StrategyQA), symbolic reasoning (Last Letter Concatenation, Coin Flip), and other logical reasoning tasks (Date Understanding, Tracking Shuffled Objects). We employ two state-of-the-art base LLMs GPT-3.5 and GPT-4 [12]. The experimental findings demonstrate significant improvements in scores compared to the zero-shot baseline across all 12 datasets. Moreover, our method not only surpasses zero-shot CoT and few-shot CoT in most arithmetic and commonsense reasoning tasks but also achieves better results when combined with zero-shot or few-shot CoT, approaching or even surpassing the performance of existing state-of-the-art methods. These results indicate the effectiveness of generating incorrect answers for individual problems to guide the reasoning process of LLMs.
@@ 2 Related Works
We provide a brief overview of the two fundamental research areas that form the basis of our work: the advent of large language models (LLMs) and prompting, and learning from negative examples.
@@
Recently, LLMs [27] like ChatGPT and GPT-4 [12] have gained significant attention. Researchers find that scaling pre-trained language models often leads to an improved model capacity on downstream tasks. These large-sized models show different behaviors from smaller models and display surprising abilities in solving a series of complex tasks.
Prompt engineering is an emerging field dedicated to the development and optimization of prompts, enabling efficient utilization of LLMs across diverse applications and research domains [1, 16]. Zero-shot prompting involves querying the LLM without any examples while few-shot prompting provides models with a few input-output examples [2]. Chain-of-thought (CoT) [21, 9] prompting enables complex reasoning capabilities through intermediate reasoning steps. Despite its success, few-shot CoT [21] needs human-labeled reasoning steps for each example, while zero-shot CoT [9] may generate incorrect reasoning steps (especially for commonsense and arithmetic reasoning). Several X-of-thought approaches [22, 23, 6, 3] extend CoT on reasoning tasks, where X can be a tree, a graph, or a program. Auto-CoT [26] improves zero-shot CoT by providing similar questions as few-shot examples for the target question. Self-consistency [19] sample multiple, diverse reasoning paths through few-shot CoT, and use the generations to select the most consistent answer. Analogical prompting [24] leverages LLMs to automatically generate relevant few-shot examples for each question. In contrast to these works, our method emphasizes eliciting self-awareness in LLMs regarding potential errors and actively avoiding them.
@@
Contrastive learning, a widely adopted technique in deep learning, aims to enhance the quality of learned representations by training models to differentiate between "positive" and "negative" samples [7]. In the LLMs area, reinforcement learning from human feedback (RLHF) [13] and direct preference optimization (DPO) [14] fine-tune LLMs with relative human judgments of response quality. Self-reflection [17, 8, 11, 25] incorporates "critic" or review steps to identify errors made by the LLM itself and improve upon them. However, it is important to note that the initial output of the LLM may not contain any errors, and there is a potential risk of the model reinforcing its own errors if it inaccurately evaluates the quality of its responses or generates invalid principles. The closest work to ours is the Contrastive CoT [4] that extends few-shot CoT by creating wrong reasoning processes from annotated correct reasoning steps. The main distinction is that the erroneous answers generated by Contrastive CoT still require human-annotated examples, and the random reordering of entities during the reasoning process may not align with the patterns of errors made by LLMs themselves. On the contrary, our approach enables LLMs to generate erroneous answers on their own, which aligns better with their intrinsic knowledge. It does not require human annotation.
@@ 3.1 Two-stage prompting
In this step we begin by transforming the input question $\mathbf{x}$ into a prompt $\mathbf{x^{\prime}}$ using a simple template “Q: [X]. A: [T]”. Here [X] represents the input slot for $\mathbf{x}$ and [T] represents a slot for a manually crafted trigger sentence $\mathbf{t}$ that would extract the reasoning process to answer the question $\mathbf{x}$. For instance, if we use “Let’s give a correct and a wrong answer.” as a trigger sentence, the prompt $\mathbf{x^{\prime}}$ would be “Q: [X]. A: Let’s give a correct and a wrong answer.”. Additional trigger examples can be found in Table 5. Prompted text $\mathbf{x^{\prime}}$ is then inputted into a LLM, which generates the subsequent sentence $\mathbf{z}$. While various decoding strategies can be employed, we have used greedy decoding throughout the paper for simplicity.
@@
In the second step, we utilize the generated sentence $\mathbf{z}$ in conjunction with the prompted sentence $\mathbf{x^{\prime}}$ to extract the ultimate answer from the LLM. To provide a more specific explanation, we combine three elements by concatenating them as "[X’] [Z] [A]". Here, [X’] represents the 1st prompt $\mathbf{x^{\prime}}$, [Z] represents the sentence $\mathbf{z}$ generated in the first step, and [A] represents a trigger sentence used to extract the answer. The prompt for this step is self-augmented, meaning that it includes the sentence $\mathbf{z}$ generated by the same LLM. During the experiment, we employed slightly different answer triggers based on the format of the answer. For instance, for multiple-choice question-answering, we use "Therefore, among A through E, the correct answer is." For math problems that require a numerical answer, we use "Therefore, the correct answer (arabic numerals) is." Please refer to Appendix A.2 for the answer trigger sentences we used in each task. Subsequently, the prompted text is inputted into the LLM to generate sentences $\mathbf{y}$ and extract the final answer.
@@ 3.2 Integrating with other prompting methods
We can easily integrate our CP with any advanced prompting methods. We name the combined method X-CP, where X can be Zero-shot-CoT, Few-shot-CoT, or any other method. X-CP also has two steps: reasoning extraction and answer extraction. For Zero-shot-CoT-CP, the only distinction is we replace the trigger sentence “Let’s give a correct and a wrong answer.” with “Let’s think step by step and give both a correct answer and a wrong answer.”. For Few-shot-CoT-CP, the distinction is that $k$ few-shot examples with reasoning steps are added before “Q: [X]. A: Let’s give a correct and a wrong answer.”, the resulting prompt $\mathbf{x^{\prime}}$ will be “Q: $X_{1}$] A: $Z_{1}$]. The answer is $Y_{1}$]. Q: $X_{2}$] A: $Z_{2}$]. The answer is $Y_{2}$]. … Q: $X_{k}$] A: $Z_{k}$]. The answer is $Y_{k}$]. Q: [X]. A: Let’s give a correct and a wrong answer.”, where $X_{i}$, $Z_{i}$ and $Y_{i}$ are the question, reasoning steps and the final answer for each example $i$.
@@ 4.1 Settings
We evaluate the effectiveness of our proposal on 12 datasets [^1] The datasets are available at https://github.com/kojima-takeshi188/zero_shot_cot/tree/main/dataset. encompassing four categories of reasoning tasks: arithmetic (SingleEq, AddSub, MultiArith, AQUA-RAT, GSM8K, SVAMP), commonsense (CommonsenseQA, StrategyQA), symbolic (Last Letter Concatenation, Coin Flip), and other logical reasoning tasks (Date Understanding, Tracking Shuffled Objects). The detailed description of each dataset can be found in [9]. We use the few-shot examples with reasoning steps provided by [21].
@@
We conducted a comprehensive comparison of our CP method with various types of prompting techniques. These include simple zero-shot methods such as Zero-shot and Zero-shot-CoT [9], Few-shot and Few-shot-CoT [21], X-of-thought approaches like Tree of Thoughts (ToT) [22], Graph of Thoughts (GoT) [23], Program-aided Language models (PAL) [6], and Program of thoughts prompting (PoT) [3]. Additionally, we compared our method with other prompting techniques such as Analogical prompting (Self-generated Exemplars) [24] and Self-consistency (SC) [19]. Furthermore, we evaluated the effectiveness of self-reflection methods, including Recursive Criticism and Improvement (RCI) [8], Self-Refine [11] and Learning Principles from Mistakes (LEAP) [25], as well as the closest related work, Contrastive CoT [4].
@@
We use GPT-4 and gpt-35-turbo (0613) as our base models (accessed between Feb 22nd–March 9th 2024). All generations are done by greedy decoding (i.e., sampling with zero temperature) as in the original CoT work [21].
@@
We follow Zero-shot-CoT [9] work and use its original implementation to pick up the final answers.
Table 1: Accuracy (in percentage) comparison of Zero-shot-CP with Zero-shot, Zero-shot-CoT and Zero-shot-CoT-CP on each dataset. The values on the left-hand side of each task represent the results obtained using GPT-4, while the values on the right-hand side represent the results obtained using gpt-35-turbo.
Arithmetic
Method SingleEq AddSub MultiArith GSM8K AQUA SVAMP
Zero-shot-CoT-CP 92.7/92.3 91.4/88.6 97.2/96.2 89.5/73.5 71.3/60.6 91.6/85.9
Common Sense Other Reasoning Tasks Symbolic Reasoning
Common Strategy Date Shuffled Last Letter Coin Flip
SenseQA QA Understand Objects (4 words) (4 times)
Zero-shot-CoT-CP 82.9/71.3 73.8/66.7 80.8/62.6 75.4/50.9 86.2/70.8 94.4/55.2

### v3

@@ Abstract
Prompting methods play a crucial role in enhancing the capabilities of pre-trained large language models (LLMs). We explore how contrastive prompting (CP) significantly improves the ability of large language models to perform complex reasoning. We demonstrate that LLMs are decent contrastive reasoners by simply adding “Let’s give a correct and a wrong answer.” before LLMs provide answers. Experiments on various large language models show that zero-shot contrastive prompting improves the performance of standard zero-shot prompting on a range of arithmetic, commonsense, and symbolic reasoning tasks without any hand-crafted few-shot examples, such as increasing the accuracy on GSM8K from $35.9\%$ to $88.8\%$ and AQUA-RAT from $41.3\%$ to $62.2\%$ with the state-of-the-art GPT-4 model. Our method not only surpasses zero-shot CoT and few-shot CoT in most arithmetic and commonsense reasoning tasks but also can seamlessly integrate with existing prompting methods, resulting in improved or comparable results when compared to state-of-the-art methods. Our code is available at the following GitHub repository: https://github.com/yao8839836/cp.
Large Language Models are Contrastive Reasoners
Liang Yao Sun Yat-sen University Shenzhen, China yaoliang3@mail.sysu.edu.cn
@@ 1 Introduction
Recent studies Zhao et al. (2023); Brown et al. (2020); OpenAI (2023) have shown that large language models (LLMs) exhibit impressive performance across a wide range of tasks. In particular, the chain-of-thought (CoT) prompting technique has demonstrated the capability of LLMs to handle complex tasks, including math problem solving, by guiding them to generate intermediate reasoning steps Wei et al. (2022b); Kojima et al. (2022); Zhang et al. (2023). These studies spotlight the significance of developing efficient techniques to direct LLMs in their reasoning processes Liu et al. (2023); Amatriain (2024); Chia et al. (2023); Yasunaga et al. (2023).
Nevertheless, the current chain-of-thought (CoT) paradigm encounters two main challenges: offering correct guidance or examples of reasoning and reducing the reliance on manual labeling. In particular, Zero-shot-CoT Kojima et al. (2022) provides general reasoning guidance by providing instructions like “Think step by step.”, but the generated reasoning steps may not be correct and adequate for tasks such as commonsense question-answering (Table 1 and 6). On the other hand, Few-shot-CoT Wei et al. (2022b) offers more detailed guidance but necessitates labeled examples of the reasoning process, which can be expensive to obtain for each task. This raises an important research question: Is it possible to generate a more accurate reasoning process without relying on human labeling?
In this work, we introduce contrastive prompting, a novel prompting approach that automatically directs the reasoning process of large language models. Our inspiration stems from how humans can learn from both their correct and incorrect actions Roediger and Finn (2009). For instance, when confronted with a math problem (as in Figure 1), people may ask "How can we prevent mistakes in each step?" By identifying the steps that are prone to mistakes on their own, they can enhance their ability to avoid mistakes and provide accurate solutions. Our idea is to prompt LLMs to emulate this reasoning process, enabling them to effectively solve new problems.
@@
Specifically, when presented with a problem to solve, we instruct LLMs to generate both correct and incorrect answers within the given context. To achieve this, we provide prompts such as "Let’s give a correct and a wrong answer." Following this, we verify and confirm the correct answer. Our proposed approach offers multiple advantages. It not only generates incorrect answers autonomously but also places a greater emphasis on ensuring the accuracy of the answers. This eliminates the need for manually labeling reasoning examples for each task and problem, effectively addressing the challenges faced by CoT.
We evaluate the proposed approach across various reasoning-intensive tasks, including arithmetic reasoning, commonsense reasoning, symbolic reasoning, and other logical reasoning tasks. We employ two state-of-the-art base LLMs GPT-3.5 and GPT-4 OpenAI (2023) and four popular open source LLMs. The experimental findings demonstrate significant improvements in scores compared to the zero-shot baseline across all datasets. Moreover, our method not only surpasses Zero-shot-CoT and Few-shot-CoT in most arithmetic and commonsense reasoning tasks but also achieves better results when combined with zero-shot or few-shot CoT, approaching or even surpassing the performance of existing state-of-the-art methods. These results indicate the effectiveness of generating incorrect answers for individual problems to guide the reasoning process of LLMs.
@@ 2 Related Works
Recently, LLMs Zhao et al. (2023) like ChatGPT and GPT-4 OpenAI (2023) have gained significant attention. Researchers find that scaling pre-trained language models often leads to an improved model capacity on downstream tasks. These large-sized models show different behaviors from smaller models and display surprising abilities in solving a series of complex tasks.
Prompt engineering is an emerging field dedicated to the development and optimization of prompts, enabling efficient utilization of LLMs across diverse applications and research domains Amatriain (2024); Sahoo et al. (2024). Zero-shot prompting involves querying the LLM without any examples while few-shot prompting provides models with a few input-output examples Brown et al. (2020). Chain-of-thought (CoT) Wei et al. (2022b); Kojima et al. (2022) prompting enables complex reasoning capabilities through intermediate reasoning steps. Despite its success, Few-shot-CoT Wei et al. (2022b) needs human-labeled reasoning steps for each example, while Zero-shot-CoT Kojima et al. (2022) may generate incorrect reasoning steps (especially for commonsense and arithmetic reasoning). Several X-of-thought approaches Yao et al. (2023a, b); Gao et al. (2023); Chen et al. (2022) extend CoT on reasoning tasks, where X can be a tree, a graph, or a program. Auto-CoT Zhang et al. (2023) improves Zero-shot-CoT by providing similar questions as few-shot examples for the target question. Self-consistency Wang et al. (2023) sample multiple, diverse reasoning paths through Few-shot-CoT, and use the generations to select the most consistent answer. Analogical prompting Yasunaga et al. (2023) leverages LLMs to automatically generate relevant few-shot examples for each question. In contrast to these works, our method emphasizes eliciting self-awareness in LLMs regarding potential errors and actively avoiding them.
@@
Contrastive learning, a widely adopted technique in deep learning, aims to enhance the quality of learned representations by training models to differentiate between "positive" and "negative" samples Jaiswal et al. (2020). In the LLMs area, reinforcement learning from human feedback (RLHF) Ouyang et al. (2022) and direct preference optimization (DPO) Rafailov et al. (2023) fine-tune LLMs with relative human judgments of response quality. Self-reflection Shinn et al. (2023); Kim et al. (2023); Madaan et al. (2023); Zhang et al. (2024) incorporates "critic" or review steps to identify errors made by the LLM itself and improve upon them. However, it is important to note that the initial output of the LLM may not contain any errors, and there is a potential risk of the model reinforcing its own errors if it inaccurately evaluates the quality of its responses or generates invalid principles. The closest work to ours is the Contrastive CoT Chia et al. (2023) that extends Few-shot-CoT by creating wrong reasoning processes from annotated correct reasoning steps. The main distinction is that the erroneous answers generated by Contrastive CoT still require human-annotated examples, and the random reordering of entities during the reasoning process may not align with the patterns of errors made by LLMs themselves. On the contrary, our approach enables LLMs to generate erroneous answers on their own, which aligns better with their intrinsic knowledge. It does not require human annotation.
@@ 3.1 Two-stage prompting
In this step we begin by transforming the input question $\mathbf{x}$ into a prompt $\mathbf{x^{\prime}}$ using a simple template “Q: [X]. A: [T]”. Here [X] represents the input slot for $\mathbf{x}$ and [T] represents a slot for a manually crafted trigger sentence $\mathbf{t}$ that would extract the reasoning process to answer the question $\mathbf{x}$. For instance, if we use “Let’s give a correct and a wrong answer.” as a trigger sentence, the prompt $\mathbf{x^{\prime}}$ would be “Q: [X]. A: Let’s give a correct and a wrong answer.”. Additional trigger examples can be found in Table 3. Prompted text $\mathbf{x^{\prime}}$ is then inputted into a LLM, which generates the subsequent sentence $\mathbf{z}$.
@@
In the second step, we utilize the generated sentence $\mathbf{z}$ in conjunction with the prompted sentence $\mathbf{x^{\prime}}$ to extract the ultimate answer from the LLM. To provide a more specific explanation, we combine three elements by concatenating them as "[X’] [Z] [A]". Here, [X’] represents the 1st prompt $\mathbf{x^{\prime}}$, [Z] represents the sentence $\mathbf{z}$ generated in the first step, and [A] represents a trigger sentence used to extract the answer. The prompt for this step is self-augmented, meaning that it includes the sentence $\mathbf{z}$ generated by the same LLM. During the experiment, we employed slightly different answer triggers based on the format of the answer. Please refer to Appendix A.2 for the answer trigger sentences we used in each task. Subsequently, the prompted text is inputted into the LLM to generate sentences $\mathbf{y}$ and extract the final answer.
@@ 3.2 Integrating with other prompting methods
We can easily integrate our CP with any advanced prompting methods. We name the combined method X-CP, where X can be Zero-shot-CoT, Few-shot-CoT, or any other method. X-CP also has two steps: reasoning extraction and answer extraction. For Zero-shot-CoT-CP, the only distinction is we replace the trigger sentence “Let’s give a correct and a wrong answer.” with “Let’s think step by step and give both a correct answer and a wrong answer.”. For Few-shot-CoT-CP, the distinction is that $k$ few-shot examples with reasoning steps are added before “Q: [X]. A: Let’s give a correct and a wrong answer.”, the resulting prompt $\mathbf{x^{\prime}}$ will be “Q: [$X_{1}$] A: [$Z_{1}$]. The answer is [$Y_{1}$]. Q: [$X_{2}$] A: [$Z_{2}$]. The answer is [$Y_{2}$]. … Q: [$X_{k}$] A: [$Z_{k}$]. The answer is [$Y_{k}$]. Q: [X]. A: Let’s give a correct and a wrong answer.”, where $X_{i}$, $Z_{i}$ and $Y_{i}$ are the question, reasoning steps and the final answer for each example $i$.
@@ 4.1 Settings
We evaluate the effectiveness of our proposal on 12 datasets [^1] The datasets are available at https://github.com/kojima-takeshi188/zero_shot_cot/tree/main/dataset. encompassing four categories of reasoning tasks: arithmetic (SingleEq, AddSub, MultiArith, AQUA-RAT, GSM8K, SVAMP), commonsense (CommonsenseQA, StrategyQA), symbolic (Last Letter Concatenation, Coin Flip), and other logical reasoning tasks (Date Understanding, Tracking Shuffled Objects). The detailed description of each dataset can be found in Kojima et al. (2022). We use the few-shot examples with reasoning steps provided by Wei et al. (2022b).
@@
We conducted a comprehensive comparison of our CP method with various types of prompting techniques. These include simple zero-shot methods such as Zero-shot and Zero-shot-CoT Kojima et al. (2022), Few-shot and Few-shot-CoT Wei et al. (2022b), X-of-thought approaches like Tree of Thoughts (ToT) Yao et al. (2023a), Graph of Thoughts (GoT) Yao et al. (2023b), Program-aided Language models (PAL) Gao et al. (2023), and Program of thoughts prompting (PoT) Chen et al. (2022). Additionally, we compared our method with other prompting techniques such as Analogical prompting (Self-generated Exemplars) Yasunaga et al. (2023) and Self-consistency (SC) Wang et al. (2023). Furthermore, we evaluated the effectiveness of self-reflection methods, including Recursive Criticism and Improvement (RCI) Kim et al. (2023), Self-Refine Madaan et al. (2023) and Learning Principles from Mistakes (LEAP) Zhang et al. (2024), as well as the closest related work, Contrastive CoT Chia et al. (2023). We also experimented with running CP using Self-consistency (SC). Specifically, we set the temperature parameter of LLMs to 0.7 and sampled 10 correct and incorrect answers. Then, we selected the answer that appeared most frequently among the 10 correct answers as the final answer.
@@
We use GPT-4 and GPT-3.5-Turbo (0613) as our base models (accessed between Feb 22nd–May 22nd 2024) for main experiments. We also tested our CP on various open LLMs: LLaMA3-8B, LLaMA3-70B Touvron et al. (2023), ChatGLM3-6B Du et al. (2022) and Qwen1.5-72B-Chat Bai et al. (2023). All generations (except experiments with Self-consistency) are done by greedy decoding (i.e., sampling with zero temperature) as in the original CoT work Wei et al. (2022b). For GPT models, we use Azure OpenAI services. For open LLMs except ChatGLM3-6B, we use LlamaAPI  [^2] https://docs.llama-api.com/quickstart. For ChatGLM3-6B, we downloaded the model and performed the inference on a Linux server with an A100 GPU.
@@
We follow Zero-shot-CoT Kojima et al. (2022) work and use its original implementation to pick up the final answers.
MultiArith GSM8K StrategyQA AddSub SVAMP CommonsenseQA
Zero-shot 60.97 14.39 65.02 82.78 69.74 71.33
Zero-shot-CoT 94.87 75.56 60.74 86.16 81.78 68.96
Zero-shot-CP 95.13 73.22 67.39 90.46 83.08 73.81
Table 1: Accuracy (in percentage) comparison of Zero-shot-CP with Zero-shot and Zero-shot-CoT on five datasets. We run all methods 5 times using GPT-3.5-Turbo and report average scores. Zero-shot-CP significantly outperforms baselines on StrategyQA, AddSub, SVAMP and CommonsenseQA based on student t-test (p < 0.05).